4x-1/3y=8
5x+1/3y=6

Answer by TimothyLamb(4379)

4x-1/3y=8
5x+1/3y=6
---
put the system of linear equations into standard form
---
4x - 1/3y = 8
5x + 1/3y = 6
---
4x - 0.333333y = 8
5x + 0.333333y = 6
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = 1.5555
y = -5.3333
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php